OpenStudy (anonymous):

Solve: sin2x cosx + cos 2x sin x = √3/2 Please help ! :)

OpenStudy (anonymous):

sinacosb+sinbcosa =sin(a+b) sin2x cosx + cos 2x sin x = √3/2 sin(2x+x)=sqrt3 /2 sin3x = sqrt 3 /2 3x= pi/3 or 2pi/3 or.... x=pi/9 or 2pi /9 or....

OpenStudy (anonymous):

that right

OpenStudy (anonymous):

how did get 3x = pi/3 ?

OpenStudy (anonymous):

how did you get 3x = pi / 3 ?

OpenStudy (anonymous):

sin30 =1/2 sin45=sqrt2 /2 sin60=sqrt 3 /2

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

inderstand now? :D

OpenStudy (anonymous):

understand*

OpenStudy (anonymous):

Ok, I understand thanks ! :)